# Improved Visual-Spatial Reasoning via R1-Zero-Like Training

### v1

@@ 1 Introduction
Following the journey of DeepSeek-R1-Zero (Guo et al., 2025), we decide to improve the visual-spatial reasoning capacities of Qwen2-VL with GRPO (Shao et al., 2024). Considering the scarcity of training data for visual-spatial question answering, we construct a video-based question answering dataset of more than 100k samples, VSI-100k, following the protocol of VSI-bench. Specifically, we leverage ScanNet (Dai et al., 2017) to get high-fidelity video scans accompanied by meticulous object-level 3D annotations, based on which (question, answer) pairs regarding spatial information can be easily crafted.
Following common practice (Guo et al., 2025; Chen et al., 2025), we define the rule-based reward function based on the alignment between the model prediction and the ground-truth answer to perform GRPO. We also include a format reward when trying to activate the CoT reasoning behaviour. The GRPO on VSI-100k turns the pretrained Qwen2-VL-2B model into the performant vsGRPO-2B within just 120 GPU hours. We observe that vsGRPO-2B outperforms the base model by 12.1%. The same pipeline also transforms the Qwen2-VL-7B model into vsGRPO-7B, achieving performance similar to that of the best open-source model with 72B parameters. During GRPO training, we have identified the necessity to keep the KL penalty (even with a small value) in the training of GRPO and observed phenomena such as reward hacking. We also compare GRPO with supervised fine-tuning (SFT) and direct preference optimization (DPO) (Rafailov et al., 2023), and confirm the superiority of GRPO in improving the spatial-visual reasoning capacities of Qwen2-VL.
@@ 3.3.3 Reward Hacking
During the training, we observed that the model sometimes finds ways to obtain high rewards that do not align with our intentions. One example is that when training with observe-mode, there are some extreme samples in the rollouts, e.g., <think> </think> <answer>xx</answer>. While this format is technically correct, it degrades to a missing observation. Observing this, we opt to incorporate a length reward function for mitigation. However, we note that some new generations just add extra <think></think> and <answer></answer> tags to exploit the length reward, which does not contribute to a meaningful thinking process. So, more reasonable reward functions should be explored.

### v2

@@ 1 Introduction
Following the journey of DeepSeek-R1-Zero (Guo et al., 2025), we decide to improve the visual-spatial reasoning capacities of Qwen2-VL with GRPO (Shao et al., 2024), sharing a similar task with the very recent work VisualThinker-R1-Zero (Zhou et al., 2025). Considering the scarcity of training data for visual-spatial question answering, we construct a video-based question answering dataset of more than 100k samples, VSI-100k, following the protocol of VSI-bench. Specifically, we leverage ScanNet (Dai et al., 2017) to get high-fidelity video scans accompanied by meticulous object-level 3D annotations, based on which (question, answer) pairs regarding spatial information can be easily crafted.
Following common practice (Guo et al., 2025; Chen et al., 2025; Zhou et al., 2025), we define the rule-based reward function based on the alignment between the model prediction and the ground-truth answer to perform GRPO. We also include a format reward when trying to activate the CoT reasoning behaviour. The GRPO on VSI-100k turns the pretrained Qwen2-VL-2B model into the performant vsGRPO-2B within just 120 GPU hours. We observe that vsGRPO-2B outperforms the base model by 12.1%. The same pipeline also transforms the Qwen2-VL-7B model into vsGRPO-7B, achieving performance similar to that of the best open-source model with 72B parameters. During GRPO training, we have identified the necessity to keep the KL penalty (even with a small value) in the training of GRPO and observed phenomena such as reward hacking. We also compare GRPO with supervised fine-tuning (SFT) and direct preference optimization (DPO) (Rafailov et al., 2023), and confirm the superiority of GRPO in improving the spatial-visual reasoning capacities of Qwen2-VL.
@@ 3.3.3 Reward Hacking
During the training, we observed that the model sometimes finds ways to obtain high rewards that do not align with our intentions. One example is that when training with observe-mode, there are some extreme samples in the rollouts, e.g., <think> </think> <answer>xx</answer>. While this format is technically correct, it degrades to a missing observation, a similar phenomenon also noted in VisualThinker-R1-Zero (Zhou et al., 2025). Observing this, we opt to incorporate a length reward function for mitigation. However, we note that some new generations just add extra <think></think> and <answer></answer> tags to exploit the length reward, which does not contribute to a meaningful thinking process. So, more reasonable reward functions should be explored.